HUBBLE SPACE TELESCOPE DAILY REPORT #1005 PERIOD COVERED: 0800 EST Daily Status Report as of 305/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) The Wide Field Camera was used to make Earth flat calibration images to be used for flat fielding of WFC observations. The observations used the narrow band filters, both with and without neutral density filters. A total of two observations were received. One observation was reported to be partially saturated, and the other was okay. 1.2 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.3 Completed PC 5023 (Calibration/WFPC Cycle 3 Polarization Calibration) The Wide Field Camera was used to make polarization calibration observations by using images of a non-polarized standard star GD319. This will allow for the absolute transmittance of the polarized optical path to be established. All images were received and displayed. No anomalies were reported. 1.4 Completed HSP/UV2 4209 (HSP Detector Sensitivity Test) The High Speed Photometer was used to make calibration observations of standard star BD75D325 to determine the instruments long term sensitivity of the IDT detectors. All observations were received and displayed. The OSS assessed the data quality, and no problems were reported. 1.5 Completed HSP/POL 4888 (Remnant Stars in Supernova Remnants - Augmentation - Cycle 3 High) The High Speed Photometer was used to make 17 ultraviolet polarimetric observations of super nova remnant SS433 which has emission lines that vary by relativistic amounts in a 162 day cycle. All seventeen observations were received and displayed. They were reported to look okay. 1.6 Completed WFC/PC 5026 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) The Wide Field Camera and the Planetary Camera were used to make internal flat calibration images. These images will be used to look for measles and daisies (i.e., different forms of contamination), and to monitor scattered light. The observations were executed as scheduled, and science data still remains of Science Tape Recorder #3. 1.7 Completed FOS/BL 4250 (Massive Stars in Starbursting Galaxies - FOS Exposures - Part 2) The Faint Object Spectrograph (Blue Side) was used to make observations of the starbursting galaxy NGC-7552 to study the physics of massive stars and the processes which drive the galaxy's formation and evolution. The OSS monitored the observations, and assessed the data quality. One recentering event was reported to have occurred. No other anomalies were reported. 1.8 Completed HSP/UV2 5084 (Visible and Ultraviolet Light Curves of Short Period RR Lyrae Type Stars) The High Speed Photometer was used to make observations of the RR Lyrae class star CY AQR (Visual Mag = 10.5). The observations were hoped to obtain high speed, high signal to noise ratio data to look for short term fluctuations in the ultraviolet. The observations were executed as scheduled, and no problems were reported. 1.9 Completed GHRS/2 2537 (Physical Conditions of Interstellar Gas Within 50 Parsecs) The High Resolution Spectrograph was used to obtain high signal to noise ratio data from the standard star HD216956 to study the warm and cold regions of the interstellar medium to analyze lines in Al, Fe, Mn, Mg, and Carbon. The observations was executed as scheduled, and no problems were reported. 1.10 Completed FOS/RD 5081 (Quasar Absorption Line Survey - cycle 3 High - Part 2) The Faint Object Spectrograph (Red Side) was used to make observations of quasar PKS0005-23 to learn more about the origin and evolution of gas along the line of sight to the target quasar. The onboard target acquisition failed, and the science observations was essentially blank. 1.11 Completed WFC 4809 (Cycle 2 : WF/PC PSF Calibration Proposal) The Planetary Camera was used to make point spread function calibration image of standard star SAO107227. The OSS monitored the observations and assessed the data quality. No problems were reported. The image was reported to look fine. 1.12 Completed Two Sets of FGS 4893 (Orbital Motions of Close Binary Stars - Cycle 3 High) Fine Guidance Sensor #3 was used to make astrometric observations of the binary star system 0537-441IN and HD80671 to obtain its orbital elements. These observations are important since most of the estimates of stellar mass are made by observing elements of these systems. The astrometry was monitored. The first observation was successful, and the second was reported to have the Stop flag set at the end of the observation. No other anomalies were reported. 1.13 Completed FGS 5057 (Extragalactic Astrometry and Astrophysics FT0-11, Part 1 of 2) Fine Guidance Sensor #3 was used to make astrometric observations of of several different extragalactic objects. These observations will be used to determine the rotation of the reference frame used by the HIPPARCOS astrometry satellite. This proposal is tied to proposal 1013. The guide star acquisition prior to the astrometry only reached Course Track, instead of the expected Fine Lock. This could degrade the quality of the astrometric observations. (See Operations Note 2.1). 1.14 Completed PC 4929 (WF/PC Cycle 2: Improved Dark Calibration) The Planetary Camera was used to make two dark calibration images. The two exposures lasted 1200 seconds each, and the camera was pre-flashed before the observations. Both observations were reported to be normal exposures. 1.15 Completed FOS/BL 4372 (Ultraviolet Spectroscopy and Spectropolarimetry of Interacting Binaries) The Faint Object Spectrograph (Blue Side) was used to make spectroscopic observations of the interacting eclipsing binary star V356-SGR. This star system has a period of 0.8 days and total eclipse lasts for about 10.6 hours. Observations were made both inside and outside the eclipse. The observations were monitored by the OSS, and the science data remains on the science tape recorder. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 17 Successful : 17 Guide star acquisition failed at 304/1527Z to obtain Fine Lock, and went to Course Track. Subsequent acquisitions were successful. Scheduled Reacquisitions : 29 Successful : 29 2.2 FHST Updates: Scheduled : 31 Successful : 31 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Five Sets of WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) 3.2 Complete FOC/96 5033 (FOC Camera, 2nd, 3rd Stage Focus - Cycle 3 ) 3.3 Complete four Sets of WFC 4381 (HST Medium - Deep Survey - Cycle 3) 3.4 Complete FOS/BL 4122 (The Stellar Content of Wolf - Rayet Galaxies (FOS Spectra)) 3.5 Complete PC 4854 (Galaxies and Clusters, WFPC GTO Augmentation, Cycle 3 High) 3.6 Complete HSP/VIS 4209 (HSP Detector Sensitivity Test) 3.7 Complete Two Sets of PC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) 3.8 Complete FOS/RD 4581 (Quasar Absorption Line Survey - Cycle 3 High) 3.9 Complete PC 4851 (Globular Cluster Cores, WFPC GTO Augmentation, Cycle 3 High) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END- HUBBLE SPACE TELESCOPE DAILY REPORT #1004 PERIOD COVERED: 0800 EST Daily Status Report as of 302/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed HSP/POL 4889 (Gravitational Lenes Part 1 - Aug. - Cycle 3 High) The High Speed Photometer was used to make nine observations of quasar QSO2345+00 which is an apparent gravitation lens system. Both photometric and polarmetric observations were made to find its characteristic signature in variability (In the time scale of weeks). All nine observations were monitored by the OSS, and the data quality was assessed. The observations were reported to be normal. 1.2 Completed PC 4308 (Snapshot Survey for Companions to Planetary Nebula Nuclei - Cycle 3 High) The Planetary Camera was used to make an observation of the central region of planetary nebula NGC-6905. It is possible that the central star in a double star. The observations was received and displayed in the OSS, and no problems were seen. 1.3 Completed FOS/RD/BL 5051 (FOS Cycle 3 Polarimetric Calibration) The Faint Object Spectrograph (Red and Blue Sides) were used to make observations of the standard star BD28D4211 to make polarametric calibration observations. The three onboard target acquisitions (all peakups) were monitored and received in real-time. The target was not bracketed by any of the peakups. 1.4 Completed WFC 4929 (WF/PC Cycle 2 : Improved Dark Calibration) The Planetary Camera was used to make two dark calibration images. The two exposures lasted 1200 seconds each, and the camera was pre-flashed before the observations. Both observations were reported to be normal exposures. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 13 Successful : 13 2.2 FHST Updates: Scheduled : 2 Successful : 2 2.3 Significant HSTARs: None 2.4 Operations Note: The SMGT #14A test terminated unexpectedly yesterday, Thursday, Oct. 27. The Primary indications are that this was caused by an incorrectly specified gyro configuration in one of the table loads generated for the test. The problem is being investigated. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.2 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.3 Complete PC 5023 (Calibration/WFPC Cycle 3 Polarization Calibration) 3.4 Complete HSP/UV2 4209 (HSP Detector Sensitivity Test) 3.5 Complete HSP/POL 4888 (Remnant Stars in Supernova Remnants - Augmentation - Cycle 3 High) 3.6 Complete WFC/PC 5026 (WF/PC cycle 3 Calibration: Rapid Internal Monitor) 3.7 Complete FOS/BL 4250 (Massive Stars in Starbursting Galaxies - FOS Exposures - Part 2) 3.8 Complete HSP/UV2 5084 (Visible and Ultraviolet Light Curves of Short Period RR Lyrae Type Stars) 3.9 Complete GHRS/2 2537 (Physical Conditions of Interstellar Gas Within 50 Parsecs) 3.10 Complete FOS/RD 5081 (Quasar Absorption Line Survey - cycle 3 High - Part 2) 3.11 Complete WFC 4809 (Cycle 2 : WF/PC PSF Calibration Proposal) 3.12 Complete Two Sets of FGS 4893 (Orbital Motions of Close Binary Stars - Cycle 3 High) 3.13 Complete FGS 5057 (Extragalactic Astrometry and Astrophysics FT0-11, Part 1 of 2) 3.14 Complete PC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 3.15 Complete FOS/BL 4372 (Ultraviolet Spectroscopy and Spectropolarimetry of Interacting Binaries) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: SMGT #14b (Session #2) will be conducted next week on Tuesday, Nov. 2, and Wednesday, Nov. 3, in two 16 hour long sessions. /DMH -END-